# Vector Quantization using Gaussian Variational Autoencoder

### v1

@@ 3.4 Grouping to Multiple Dimensions
TR is the least flexible strategy and must be used during the training of the Gaussian VAE. Specifically, building on the PT quantization approach, we can relax the TDC training target by considering the relationship between $\sum_{l=i}^{i+m}D_{KL}(q(Z_{l}|X)||\mathcal{N}(0,1))$ and $\log_{2}K\pm 0.5$ bits. In terms of performance, PQ does not affect reconstruction at all. PT provides slight improvements in reconstruction, while TR significantly enhances reconstruction performance (see Table 13).
@@ 3.5 Improving TokenBridge with Target Divergence Constraint
TokenBridge (wang2025bridging) also convert a pre-trained Gaussian VAE into a VQ-VAE. It adopts the Post Training Quantization (PTQ) technique from model compression and proposes to treat latent as model parameters to discretize. It uses a fixed codebook composed of $2^{K}$ centroids of a Gaussian distribution. It then quantizes the posterior sample by finding the closest centroid. However, TokenBridge directly quantizes a vanilla Gaussian VAE without limiting the KL divergence of each dimension. This leads to suboptimal rate-distortion performance. In fact, we can also improve the performance of TokenBridge using TDC. Specifically, the quantization centers of TokenBridge are the equal-probability partition centers of the $\mathcal{N}(0,1)$ distribution, which can be seen as a special case of GQ with an evenly distributed codebook $c_{1:K}$. From this perspective, the number of PTQ bits should also match the bits-back coding bitrate of the Gaussian VAE, and TDC can therefore enhance the performance of TokenBridge. As we demonstrate in Table 3, TDC indeed improves TokenBridge performance by a large margin.
3.6 Relationship with Reverse Channel Coding
GQ is closely related to reverse channel coding, which aims to simulate a distribution $q$ using samples from a distribution $p$ (harsha2007communication; li2018strong; havasi2018minimal; Flamich2020CompressingIB; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating). For example, Minimal Random Coding (MRC) (havasi2018minimal), when applied to a Gaussian VAE, samples from the categorical distribution with logits given by the likelihood difference:
$\hat{z}_{i}\sim\hat{q}(c_{1:K})\propto e^{\log q(Z_{i}=c_{j}|X)-\log\mathcal{N}(c_{j}|0,1)}.$ (12)
The key difference between MRC and GQ is that MRC and its variants (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating) simulate a distribution through stochastic sampling, whereas a VQ-VAE requires deterministic quantization. For one dimension quantization, the bias bound of MRC derived from Chatterjee2015TheSS can not be achieved as VQ-VAE does not allow stochastic encoding. On the other hand, our achievability and converse bound is compatible with deterministic quantization. In terms of quantization error, GQ outperforms MRC by definition (Eq. 3). Besides, GQ without grouping (m=1) can be implemented by bisect search with better asymptotic complexity (See Appendix D.11).
Additionally, TDC is closely related to the MIRACLE heuristic and IsoKL parametrization of Gaussian VAE (Havasi2018MinimalRC; Flamich2022FastRE; Lin2023MinimalRC). More specifically, MIRACLE also proposes adjusting $\lambda$ during VAE training. However, MIRACLE only maintains a single $\lambda$, making it less effective for controlling the mean of $D_{KL}$ but less effective for constraining its minimum and maximum values. On the other hand, IsoKL imposes strict control on $D_{KL}$ by directly solving for $\sigma$ given $\mu$ using the Lambert W function (corless1996lambert; brezinski1996extrapolation). However, IsoKL suffers from numerical issues and leads to suboptimal performance.
Figure 2: Qualitative results on ImageNet dataset and 0.25 bpp. Our GQ has most visually pleasing reconstruction result.
@@ 4.1 Experimental Setup
Models and Baselines For image reconstruction, we select two representative autoencoder architectures: UNet from Stable Diffusion 3 (Esser2024ScalingRF), and ViT from the BSQ (zhao2024image). For VQ-VAE baselines, we include vanilla VQGAN (van2017neural) and several representative variants, including FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image). Besides, we compare our approach to pre-trained VQ-VAE of VQGAN-Taming (esser2021taming), VQGAN-SD (rombach2022high) Llama-Gen (sun2024autoregressive), FlowMo (Sargent2025FlowTT), BSQ (zhao2024image) and other conversion approaches such as TokenBridge and ReVQ . Additionally, we demonstrate that our TDC technique can improve TokenBridge (wang2025bridging), a previous training-free approach for converting Gaussian VAEs into VQ-VAE. For image generation, we employ the Llama transformer (Touvron2023Llama2O; Shi2024ScalableIT).
Datasets, Bitrates, and Metrics For the datasets, we use the ImageNet (Deng2009ImageNetAL) training split for training, and both the ImageNet and COCO (Lin2014MicrosoftCC) validation split for testing. For reconstruction and generation experiments, all images are resized to $256\times 256$ and $128\times 128$ respectively. In terms of bitrates, we evaluate the image reconstruction performance of all models using a codebook size of $2^{14}-2^{18}$ and token numbers of $1024$, $2048$, and $4096$, which correspond to bpp (bits-per-pixel) values of 0.221.000.22-1.00. This extends the BSQ evaluation beyond 0.250.500.25-0.50 bpp. For metrics, we adopt Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang2018TheUE), Structural Similarity Index Measure (SSIM) (Wang2004ImageQA), and reconstruction Fréchet Inception Distance (rFID) (Heusel2017GANsTB) for image reconstruction; generation Fréchet Inception Distance (gFID) and Inception Score (IS) (Salimans2016ImprovedTF) are used for image generation. For further details, see Appendix C.
Table 1: Quantitive results on ImageNet dataset. Our GQ outperforms other VQ-VAEs on both UNet and ViT architecture, across 0.25-1.00 bpp. Bold: best.
Method
bpp
(# of tokens)
UNet based ViT based
[…]
VQGAN
0.25
[…]
26.51 0.125 0.748 5.714 25.39 0.103 0.740 3.518
FSQ 26.34 0.075 0.756 1.125 25.87 0.109 0.751 3.856
LFQ 24.74 0.164 0.722 16.337 24.81 0.143 0.725 15.716
BSQ 25.62 0.086 0.754 1.080 26.52 0.083 0.793 1.649
GQ (Ours) 27.61 0.059 0.807 0.529 27.88 0.061 0.823 0.932
VQGAN
0.50
($^{\textup{16}}\times2$048)
29.21 0.052 0.831 1.600 27.86 0.062 0.823 1.228
FSQ 29.29 0.047 0.845 0.871 28.83 0.055 0.842 1.067
LFQ 26.90 0.107 0.800 8.035 27.87 0.068 0.829 2.444
BSQ 27.88 0.059 0.836 0.788 28.44 0.051 0.852 0.700
GQ (Ours) 30.17 0.039 0.875 0.492 30.42 0.037 0.882 0.592
VQGAN
1.00
[…]
32.06 0.026 0.896 0.580 31.32 0.032 0.899 0.716
FSQ 32.38 0.025 0.905 0.636 31.58 0.026 0.905 0.544
LFQ 28.31 0.074 0.840 3.617 26.67 0.105 0.790 8.288
BSQ 30.50 0.032 0.900 0.346 31.60 0.027 0.914 0.379
GQ (Ours) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
Table 2: Quantitive results on ImageNet dataset. Our GQ outperforms previous pre-trained VQ-VAEs with less training. Bold: best, : from paper, -: not available.
Method
bpp
(# of tokens)
[…]
ImageNet
@@
VQGAN-Taming
0.22
@@
23.38 - - 1.190 (OpenImages) 67
VQGAN-SD - - - 1.140 (OpenImages) 83
Llama-Gen-32 24.44 0.064 0.768 0.590 40 70
FlowMo-Hi 24.93 0.073 0.785 0.560 300 945
GQ (Ours) 25.31 0.064 0.762 0.491 40 82
BSQ
0.28
@@
Table 3: Quantitive results of improving TokenBridge based on UNet architecture. Our TDC improves TokenBridge significantly.
Method
bpp
(# of tokens)
ImageNet validation COCO validation
@@
4.2 Main Results
Image Reconstruction In Table 1 and Table 15, we compare our GQ method to other quantization approaches across the 0.250.251.001.00 bpp range. The results show that, in terms of reconstruction metrics such as PSNR, LPIPS, SSIM, and rFID, our GQ approach achieves state-of-the-art performance in most cases. The advantage of GQ is consistent across both UNet and ViT model architectures, as well as for both the ImageNet and COCO datasets. Visually, in Figure 2, it is shown that our GQ also produces pleasing reconstruction and preserves a lot more details in the source image. Besides, in Table. 2, we show that our GQ achieves competitive performance compared with several pre-trained models such as FlowMo, with less training epochs. Additionally, in Table. 11, we show that our GQ outperforms previous methods to discretize Gaussian VAE, including TokenBridge and ReVQ.
Improving TokenBridge In Table 3, we compare TokenBridge (wang2025bridging) applied to a vanilla Gaussian VAE and a TDC constrainted Gaussian VAE. The results show that the quantization error of TokenBridge is quite large when applied to a vanilla Gaussian VAE. In contrast, TDC significantly reduces the quantization error.
Image Generation In Table 4, we evaluate the performance of GQ in terms of image generation. It is shown that compared with VQGAN, FSQ, LFQ and BSQ, our GQ has higher codebook usage and codebook entropy. In terms of generation FID and IS, our GQ is comparable to FSQ and better than other methods. Additionally, we train a DiT (Peebles2022ScalableDM) with same model architecture and same training setting, using the Gaussian VAE with and without TDC. It is shown that for limited computation, auto-regressive generation is more efficient than diffusion generation, in terms of both FID and IS. This result shows that the conversion from Gaussian VAE to VQ-VAE facilities auto-regressive generation, which improves the efficiency of image generation.
Complexity Compared with Gaussian VAE, the overhead of GQ is negligible (See Appendix D.10).
Table 4: Quantitive results on class conditional image generation of ImageNet dataset. Our GQ achieves best codebook usage and competitive generation performance. Bold: best.
@@ 4.2 Main Results
Diffusion
Gaussian VAE w/o TDC - - 8.35 202.19
Gaussian VAE w/ TDC - - 8.47 205.94
Auto-regressive
VQGAN 16.4% 4.36 8.01 151.40
FSQ 94.3% 14.74 7.33 224.88
LFQ 24.9% 9.65 7.73 142.09
BSQ 99.8% 14.93 7.82 221.64
TokenBridge 94.6% 14.94 7.82 198.24
GQ (Ours) 100.0% 15.17 7.67 230.79
4.3 Ablation Studies
Effectiveness of Pre-trained Gaussian VAE It is possible to train a vanilla VQ-VAE (van2017neural) using the same codebook as GQ, which is equivalent to a VQ-VAE with a fixed, Gaussian noise codebook. It is also possible to directly train the Gaussian VAE neural network with the GQ target in Eq. 11 using Gumbel-Softmax (jang2016categorical; maddison2016concrete). However, as shown in Table 5, both methods do not converge well. Furthermore, fine-tuning GQ after initializing it with a pre-trained Gaussian VAE also has only marginal effect on performance. These results indicate that GQ’s conversion from a pre-trained Gaussian VAE is necessary and sufficient.
Table 5: The effect of pre-trained Gaussian VAE. Converting GQ from pre-trained Gaussian VAE is better than training GQ from scratch.
@@ 4.3 Ablation Studies
GQ (from scratch) Vanilla VQ-VAE 1.00 8.50 0.763 0.156 360.597
GQ (from scratch) Gumbel-softmax Eq. 11 29.65 0.044 0.866 0.928
GQ (finetune from Gaussian VAE) Gumbel-softmax Eq. 11 32.45 0.022 0.905 0.264
GQ (convert from Gaussian VAE) no 32.47 0.023 0.907 0.327
Effectiveness and Alternatives of TDC To demonstrate the necessity of TDC in Eq. 9, we train a vanilla Gaussian VAE without TDC. As shown in Table 6, the mean $D_{KL(2)}$ of the vanilla Gaussian VAE is close to that of the Gaussian VAE with TDC (3.99 vs. 4.26 bits). However, the range of $D_{KL(2)}$ is much wider for the vanilla model (0.26–27.29 vs. 2.93–5.63 bits). Although the reconstruction performance of the two Gaussian VAEs is very similar (PSNR: 32.73 vs. 32.61 dB, rFID: 0.490 vs. 0.460), GQ with TDC outperforms GQ without TDC by a large margin (PSNR: 31.25 vs. 26.43 dB, rFID: 0.372 vs. 0.978). This demonstrates that TDC is necessary.
Alternatives to TDC are the MIRACLE heuristic (Havasi2018MinimalRC) and IsoKL (Flamich2022FastRE). MIRACLE is not that effective in terms of controlling the range is outperformed by TDC (PSNR 29.48 vs. 32.11 dB). On the other hand, IsoKL imposes a stricter constraint by requiring that $D_{KL}$ is exactly the same across all dimensions. Iso-KL enforece the constraint well but has inferior performance (PSNR 30.45 vs. 32.11 dB). This is because IsoKL is not numerically stable and it discards the solution with $\sigma^{2}_{i}>1$. In Appendix B, we propose a numerically stable version of Mean-KL (Lin2023MinimalRC), which is a IsoKL that supports grouping ($m>1$). However, it does not work well for ViT based models.
Table 6: The effect of adding constraints to Gaussian VAE. GQ is effective only on Gaussian VAE trained with TDC constraint.
[…]
Gaussian VAE no 3.99, 0.26-27.29 - 1.00 32.73 0.022 0.910 0.490
GQ - - 4 1.00 26.43 0.834 0.054 0.978
@@
Alternatives to GQ There are several stochastic alternatives to GQ, such as MRC, ORC, and A coding (havasi2018minimal; Theis2021AlgorithmsFT; Flamich2022FastRE; he2024accelerating). In Table 7, we compare these methods in terms of reconstruction quality. When applied to a TDC-constrained Gaussian VAE, GQ achieves the best PSNR, SSIM, and rFID. Besides, when grouping $m=1$, GQ can be implemented using bisect search. This means that GQ is asymptotically faster than MRC, ORC and A coding (See Appendix D.11).
Table 7: Comparison between MRC methods and GQ on ImageNet dataset. GQ has better reconstruction quality and can be implemented asymptotically faster when grouping $m=1$.

### v2

@@ 3.4 Supporting Multi-dimensional Codebook
Now we have the conversion method for multi-dimensional codebook, we consider how to train a Gaussian VAE for multi-dimensional codebook. Specifically, we modify the TDC training target to account for the relationship between the sum of $m$ bitrates $\sum_{k=i}^{i+m}R_{k}$ and $\log_{2}K\pm\alpha$. The detailed TDC update rule is:
$\mathcal{L}_{TDC}^{m}=\sum_{j=1}^{d//m}\lambda_{j}\sum_{k=jm}^{jm+m}R_{k}+\mathbb{E}[\Delta(X,g(z))],$
$\textup{where }\lambda_{i}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&\sum_{k=jm}^{jm+m}R_{k}<\log_{2}K-\alpha,\\ \lambda_{\textrm{mean}},&\sum_{k=jm}^{jm+m}R_{k}\in[\log_{2}K\pm\alpha],\\ \lambda_{\textrm{max}},&\sum_{k=jm}^{jm+m}R_{k}>\log_{2}K+\alpha.\end{array}\right$. (14)
[…]
Table 1: Quantitative results on the ImageNet dataset, comparing different VQ methods using the same model architecture. Our GQ outperforms other VQ methods on both UNet and ViT architectures. Bold: best. Gray: continuous Gaussian VAE.
Method bpp ($K\times N$) UNet based ViT based
[…]
Gaussian VAE $\approx$ 0.25 (-) 28.60 0.047 0.849 0.556 28.95 0.045 0.851 0.672
Gaussian VAE (w/ TDC) 28.05 0.053 0.829 0.535 28.44 0.054 0.837 0.793
[…]
FSQ (Mentzer et al., 2023) 26.34 0.075 0.756 1.125 25.87 0.109 0.751 3.856
LFQ (Yu et al., 2023) 24.74 0.164 0.722 16.337 24.81 0.143 0.725 15.71
BSQ (Zhao et al., 2024) 25.62 0.086 0.754 1.080 26.52 0.083 0.793 1.649
GQ (Ours) 27.61 0.059 0.807 0.529 27.88 0.061 0.823 0.932
Gaussian VAE $\approx$ 0.50 (-) 31.18 0.028 0.889 0.389 31.64 0.030 0.900 0.867
Gaussian VAE (w/ TDC) 30.84 0.035 0.884 0.549 31.18 0.033 0.896 0.757
VQGAN (Esser et al., 2021) 0.50 ($^{\textup{16}}\times2$048) 29.21 0.052 0.831 1.600 27.86 0.062 0.823 1.228
FSQ (Mentzer et al., 2023) 29.29 0.047 0.845 0.871 28.83 0.055 0.842 1.067
LFQ (Yu et al., 2023) 26.90 0.107 0.800 8.035 27.87 0.068 0.829 2.444
BSQ (Zhao et al., 2024) 27.88 0.059 0.836 0.788 28.44 0.051 0.852 0.700
GQ (Ours) 30.17 0.039 0.875 0.492 30.42 0.037 0.882 0.592
Gaussian VAE $\approx$ 1.00 (-) 32.73 0.022 0.910 0.490 32.28 0.023 0.910 0.469
Gaussian VAE (w/ TDC) 32.75 0.021 0.910 0.294 32.05 0.023 0.909 0.397
[…]
FSQ (Mentzer et al., 2023) 32.38 0.025 0.905 0.636 31.58 0.026 0.905 0.544
LFQ (Yu et al., 2023) 28.31 0.074 0.840 3.617 26.67 0.105 0.790 8.288
BSQ (Zhao et al., 2024) 30.50 0.032 0.900 0.346 31.60 0.027 0.914 0.379
GQ (Ours) 32.47 0.023 0.907 0.322 31.71 0.024 0.903 0.349
3.5 Practical Guidelines
To summarize, GQ has the same interface as a VQ-VAE and constructs a VQ-VAE with codebook size $K$, codebook dimension $m$, and number of tokens $N$ with two steps:
  • First, GQ trains a Gaussian VAE with latent dimension $d=mN$, using the TDC constraint with target $\log K$ as in Eq. 14.
  • Next, GQ converts this constrained Gaussian VAE into a VQ-VAE using Eq. 10. The codebook $c_{1:K}$ is fixed once generated.
@@ 3.6 Improving TokenBridge with Target Divergence Constraint
TokenBridge (Wang et al., 2025) also converts a pre-trained Gaussian VAE into a VQ-VAE. It adopts the Post-Training Quantization (PTQ) technique from model compression, treating the latent as model parameters for discretization. TokenBridge uses a fixed codebook of $2^{K}$ centroids from a Gaussian distribution and quantizes the posterior by selecting the closest centroid. However, it directly quantizes a vanilla Gaussian VAE without constraining the KL divergence of each dimension, leading to suboptimal rate-distortion performance.
We can improve TokenBridge using TDC. Its quantization centers correspond to equal-probability partition centers of $\mathcal{N}(0,1)$, which is a special case of GQ with an evenly distributed codebook $c_{1:K}$. Therefore, the number of PTQ bits should also match $R_{i}$, and TDC can enhance TokenBridge performance (see Table 3).
@@ 4.1 Experimental Setup
Models and Baselines For image reconstruction, we use two representative autoencoder architectures: UNet from Stable Diffusion 3 (Esser et al., 2024), and ViT from BSQ (Zhao et al., 2024). For VQ baselines, we train vanilla VQGAN (Esser et al., 2021) and several variants, including FSQ (Mentzer et al., 2023), LFQ (Yu et al., 2023), and BSQ (Zhao et al., 2024) on the same model architecture. Additionally, we compare to pre-trained VQ-VAEs such as VQGAN-Taming (Esser et al., 2021), VQGAN-SD (Rombach et al., 2022), Llama-Gen (Sun et al., 2024a), FlowMo (Sargent et al., 2025), BSQ (Zhao et al., 2024), OpenMagViT-V2 (Luo et al., 2024), and conversion methods like TokenBridge and ReVQ. We further show that our TDC technique improves TokenBridge (Wang et al., 2025), a previous training-free approach for converting Gaussian VAEs into VQ-VAEs. For image generation, we employ the Llama transformer (Touvron et al., 2023; Shi et al., 2024a).
Datasets, Bitrates, and Metrics We use the ImageNet (Deng et al., 2009) training split for training, and the ImageNet and COCO (Lin et al., 2014) validation splits for testing. For reconstruction and generation experiments, images are resized to $256\times 256$ and $128\times 128$, respectively. For bitrates, we evaluate reconstruction performance using codebook sizes K from $2^{14}$ to $2^{18}$ and number of token N of $256$, $1024$, $2048$, and $4096$. We compute bits-per-pixel (bpp) as $\log_{2}K\times N/H\times W$, where $H,W$ are image height and width. In our case, we cover bpp values of 0.070.071.001.00. We fix channel dimension to 16 following original Stable Diffusion 3 UNet. For bpp $0.50,1.00$ we adopt codebook dimension $4,8$ respectively. For bpp <0.50<0.50, we adopt codebook dimension 16. For either UNet or ViT architecture, we follow ReVQ (Zhang et al., 2025) to put codebook dimension on channel dimension. Metrics include Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and reconstruction Fréchet Inception Distance (rFID) (Heusel et al., 2017) for image reconstruction; and generation Fréchet Inception Distance (gFID) and Inception Score (IS) (Salimans et al., 2016) for image generation. Further details are provided in Appendix C.
Table 2: Quantitative results on the ImageNet dataset, compared with pre-trained VQ-VAEs. Our GQ outperforms prior pre-trained VQ-VAEs while requiring less training. Bold: best, : from paper, -: not available.
@@
OpenMagViT-V2 (Luo et al., 2024) 0.07 ($^{\textup{14}}\times2$56) 21.63 0.111 0.640 1.17 300 170
ReVQ-256T (Zhang et al., 2025) 21.96 0.121 0.640 2.05 (Mixed dataset as DC-AE) 83
GQ (Ours) 22.30 0.116 0.642 1.04 40 87
@@
VQGAN-SD (Rombach et al., 2022) - - - 1.140 (OpenImages) 83
Llama-Gen-32 (Sun et al., 2024a) 24.44 0.064 0.768 0.590 40 70
FlowMo-Hi (Sargent et al., 2025) 24.93 0.073 0.785 0.560 300 945
GQ (Ours) 25.31 0.064 0.762 0.491 40 87
@@
Table 3: Quantitative results of improving TokenBridge on the UNet architecture. Applying TDC significantly enhances the reconstruction quality of TokenBridge.
Method bpp ($K\times N$) ImageNet validation COCO validation
@@
Table 4: Comparison between GQ conversion and its alternatives on the ImageNet dataset. GQ achieves better reconstruction quality and can be implemented asymptotically faster when the codebook dimension $m=1$. Here, $D_{\infty}(\cdot||\cdot)$ denotes the Rényi $\infty$-divergence.
@@
Gaussian VAE (w/ TDC) O(1)O(1)/$O(1)$ 32.75 0.023 0.906 0.460
MRC (original) $O(2^{R_{i}})$/$O(1)$ 32.09 0.023 0.906 0.425
ORC $O(2^{R_{i}})$/$O(1)$ 32.09 0.023 0.906 0.419
GQ (Ours) $O(R_{i})$/$O(1)$ 32.11 0.023 0.907 0.414
Table 5: Effect of TDC and alternatives. GQ is effective only when applied to a Gaussian VAE trained with the TDC constraint.
@@
Gaussian VAE None 3.99, 0.26-27.29 - 1.00 32.73 0.022 0.910 0.490
GQ - - 4 1.00 26.43 0.054 0.834 0.978
@@
4.2 Main Results
Image Reconstruction Tables 1 and 15 compare GQ with other quantization methods with same model architecture. Across reconstruction metrics such as PSNR, LPIPS, SSIM, and rFID, GQ achieves state-of-the-art performance in most cases. Its advantage is consistent across UNet and ViT architectures, as well as on the ImageNet and COCO datasets. Moreover, the quantization brought by GQ is marginal compared to Gaussian VAE. Figure 2 shows that GQ produces visually pleasing reconstructions, preserving substantially more details from the source image. Moreover, Table 2 shows that GQ achieves competitive performance relative to several pre-trained models, such as FlowMo, while requiring fewer training epochs and smaller or comparable model size. Finally, GQ outperforms prior Gaussian VAE discretization methods, including TokenBridge and ReVQ.
Improving TokenBridge Table 3 compares TokenBridge (Wang et al., 2025) applied to a vanilla Gaussian VAE versus a TDC-constrained Gaussian VAE. The results show that TokenBridge incurs substantial quantization error on a vanilla Gaussian VAE, whereas applying TDC significantly reduces this error.
Table 6: Quantitative results for class-conditional image generation on the ImageNet dataset. GQ achieves the highest codebook utilization while maintaining competitive generation performance. Bold: best. Underline: second best.